# Split Unlearning

### v3

@@ 4.2 Masked Label Sharing in SplitWiper+
Masked label sharing. In SplitWiper+, after freezing the pre-trained model, clients apply an expansion-and-masking scheme before sharing their data with the server. In contrast, SplitWiper involves clients directly sharing the obtained ${Y_{q},V_{q}^{k}}$ with the server after freezing the pre-trained model, without the additional masking step. By transforming $\{Y_{q},V_{q}^{k}\}$ into $\{Y_{e}^{*},{V_{e}^{*}}^{k}\}$, clients transmit the processed masked labels and their corresponding DP-protected intermediate values to the server, without revealing the original information.
@@
Consider that $\mathbb{V}_{exp}=\{{V_{1}^{*}},{V_{2}^{*}},\cdots,{V_{E}^{*}}\}$ is expanded by $\mathbb{V}=\{{V}_{1},{V}_{2},\cdots,{V}_{Q}\}$ using a given method $\mathcal{G}_{V}$, where $\forall{V}_{q}\in\mathbb{V}$, $\gamma_{q}$ copies with $\varepsilon$-DP-based noise are generated as $\{{V_{a_{1}}^{*}},\cdots,{V_{a_{\gamma_{s}}}^{*}}\}\in\mathbb{V}_{exp}$. The probability that any two elements in $\mathbb{V}_{exp}$ correspond to the same element in $\mathbb{V}$ satisfies $2\varepsilon$-DP.
@@
We introduce an additional label demasking step on the client side when a client needs to utilize the complete model, i.e., $Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s})$. This step converts the predicted masked labels from the server back into their original, unmasked forms to fulfill the task at hand. Given that $\mathcal{G}_{Y}$ is a one-to-many expansion method, its inverse, $\mathcal{G}_{Y}^{-1}$, is a many-to-one function as defined:
@@
Each masked label corresponds directly to a real label, and since the mapping $\mathcal{G}_{Y}$ is known to all clients, they can effortlessly demask the predicted results received from the server by applying the inverse mapping $\mathcal{G}_{Y}^{-1}$. This demasking process is not merely a reversal but an integral part of maintaining the integrity and applicability of the model’s predictions. The demasked labels are then used to compute the performance $\mathcal{T}(Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s}),D_{o}^{k})$, ensuring that the model evaluation is based on accurate and relevant data, reflective of true model performance in operational settings.
Label protection in the unlearning process. When the initiating client-$k$ requests unlearning, only client-$k$ participates in the subsequent update process. Other clients do not receive any additional updated information. The server, upon receiving the updated masked labels and corresponding DP-protected intermediate values shared by client-$k$, gains insight into the aspects of the unlearning process. However, since the labels are masked, the server cannot accurately infer the quantity or category of the real labels involved in the unlearning from the changes in the number of masked labels. Thus, we deem that the privacy associated with the unlearning labels remains protected, effectively safeguarding sensitive information from inadvertent disclosure.
@@
    Training on private datasets. When training on private datasets, an external attacker, even with partial access to a victim’s data, cannot construct a comprehensive shadow dataset. Although [21] assumes the attacker knows all labels in the victim’s dataset, this is unrealistic unless using a well-known public dataset. The need to balance positive and negative cases and diversify shadow models is crucial for generalization in the attack model. Without full label knowledge, inferring the complete input-output mapping becomes difficult, limiting the attacker’s ability to replicate the training environment and conduct effective inference.
@@
Compared to SplitWiper, the label expansion strategy in SplitWiper+ introduces additional computational overhead and increases the size of the data transmitted from clients to the server, thereby increasing communication costs. Thanks to the one-way-one-off propagation scheme, these additional overheads are incurred only once, preserving the significant advantages over existing SL methods. Table 1 presents a comparison of the complexities among SplitWiper, SplitWiper+, and existing Vanilla and U-shaped SL. A detailed analysis of complexities can be found in Appendix B.
Table 1: Client-side Privacy and Overhead in SL
Framework Label Privacy Computational Communication
SplitWiper Low $O(\mathcal{H}_{K})$ $O(\mathcal{I}_{K}+\mathcal{I}_{Y})$
@@ 5.1 Experimental Settings
SplitWiper and SplitWiper+ are versatile, accommodating a variety of unlearning scenarios including class-, client-, and sample-level tasks. The experiments specifically focus on class-level unlearning tasks for clarity of presentation [24], where $\mathbb{C}_{u}$ selects a subset of label, denoted as $\mathbb{Y}_{u}$ for unlearning, aiming to effectively erase all related information without affecting other clients $\mathbb{C}_{o}$, or other labels $\mathbb{Y}_{o}$. We set $\lvert\mathbb{Y}_{u}\rvert=1$, i.e., only a specific label is required for unlearning. The unlearned label set, $\mathbb{Y}_{u}$, is exclusively possessed by $\mathbb{C}_{u}$, aligned with the non-IID data distribution setting and enabling a clearer demonstration of unlearning effectiveness.
@@
    Clients with different tasks. Clients utilize different datasets for separate tasks, with completely distinct labels among the clients.
@@
For experiments with clients performing the same task, we use CIFAR-10 [25] and CIFAR-100 [26] datasets, which are widely utilized in the fields of computer vision and neural networks. CIFAR-10 is a collection of $60,000$ $32\times 32$ color images across $10$ classes, and CIFAR-100 includes 100 classes with $600$ images each. In CIFAR-10, we designate the label $Y_{u}$ as “bird”, and in CIFAR-100, as “aquarium-fish”. As for clients with different tasks, $3$ clients use CIFAR-10 and $2$ clients use MNIST [27] that includes a dataset of $60,000$ $28\times 28$ single-channel handwritten digit images. In this setup, “bird” from CIFAR-10 remained as $Y_{u}$. These diverse experimental setups help illuminate the strengths and challenges of implementing SplitWiper across different scenarios and datasets, reflecting its effectiveness in managing complex unlearning demands in non-IID distributed SL environments.
@@
In scenarios where clients train on the same task, each client’s model, $\mathcal{F}_{o}^{k},\forall k$, adheres to a uniform structure with the same input and output sizes across all clients. Conversely, in scenarios with clients engaged in different tasks, the input size of each $\mathcal{F}_{o}^{k}$ is tailored based on the specific training dataset $D_{o}^{k}$ used locally. However, the output size remains consistent to meet the input requirements of the server model $\mathcal{F}_{o}^{s}$. This design ensures that despite the diverse data types and tasks, all client models can interface correctly with the centralized server model within the SL framework.
@@
Benchmarks and experimental groups. We employ existing Vanilla and U-shaped SL [4] as benchmarks to evaluate the performance of SplitWiper. To ensure a fair comparison, we initialize Vanilla and U-shaped SL with the pre-trained results from SplitWiper. Then, both the client and server in these existing frameworks are set to undergo the same $M$ epochs as specified in SplitWiper to train in the learning task or re-train when asking for unlearning. By aligning the initial conditions and training durations across all frameworks, we assess the efficiency and effectiveness of SplitWiper relative to the established benchmarks.
@@
To assess the impact of different label expansion strategies in SplitWiper and SplitWiper+, we categorize the experiments based on various expansion factors. Specifically, four groups are created: three with uniform label expansions of $\gamma=1$, $2$, or $3$. When $\gamma=1$ (SplitWiper), labels are shuffled and re-indexed, but their actual count remains unmasked, similar to Vanilla SL where real labels are shared. For $\gamma=2$ or $\gamma=3$ (SplitWiper+), clients expand labels to $\gamma$ instances and apply $\varepsilon$-DP to protect the corresponding intermediate values, with $\varepsilon=0.1$ and $0.2$. We also experiment with a scenario denoted as $\gamma=\Gamma_{p}$ where $\gamma_{q}$ for each $Y_{q}$ is randomly selected within a given range $\Gamma_{p}=[1,3]$. This method allows us to examine how variability in the expansion factor affects label privacy and system performance.
Performance metrics. We use training and transmission times as primary indicators of computational and communication overheads for different SL frameworks to highlight the superiority of the proposed one-way-one-off propagation. We focus on client-side overhead due to the comparable server-side computational costs across frameworks. In SplitWiper and SplitWiper+, the server-side complexity is $O(M\cdot\mathcal{H}_{S})$, where $M$ is the number of training epochs and $\mathcal{H}_{S}$ is the complexity of a single server training iteration. Similarly, Vanilla SL has a server-side complexity of $O(M\cdot\mathcal{H}_{S})$, while U-shaped SL reduces this to $O(M\cdot(\mathcal{H}_{S}-\mathcal{H}_{out}))$ due to the output layer’s location in the client model. We consider a bandwidth of $100$ Mbps to simulate transmission costs between clients and the server. We evaluate the unlearning effectiveness of each framework by comparing the label accuracy during both the learning and unlearning phases. Our experiments run on NVIDIA PCIe A100, 2 $\times$ 40GB, and we use Python 3.7.16 and PyTorch 2.3.1 to build and train the proposed framework.
@@ 5.2 Unlearning Efficiency and Overhead
The results highlight the efficacy of SplitWiper in achieving G1 (independent unlearning). Training and transmission times in SplitWiper are significantly lower than those in Vanilla and U-shaped SL across all data distributions and scenarios. Crucially, during unlearning, only the requesting clients, $\mathbb{C}_{u}$, incur additional costs, while other clients, $\mathbb{C}_{o}$, remain unaffected, as the server uses cached data from the learning phase to handle unlearning, minimizing overall client-side expenses and disruptions. The time overhead for the label expansion strategy discussion between clients (cf. Algorithm 3) is negligible, measured in milliseconds, and does not significantly impact our analysis of training and transmission times.
In SplitWiper and SplitWiper+, clients only need to train the pre-trained model over local data once. They then mask labels according to the label expansion strategy and apply DP protection to the corresponding intermediate values before transmitting them to the server in a one-way propagation manner. Subsequent training phases focus exclusively on the server, disconnecting from client-side computations. Therefore, the training and transmission times for clients are fixed values dependent on $\gamma$ and remain unaffected by the number of training epochs $M$. In contrast, in Vanilla and U-shaped SL, where each epoch involves client participation, the computational and transmission times increase linearly with $M$, as listed in Table 1 (see Appendix B).
@@
To intuitively demonstrate the impact of the label expansion factor $\gamma$ on client-side overhead in SplitWiper+, we plot the average training and transmission times for all clients across different $\gamma$ values in the learning scenarios in Fig. 4(a) and Fig. 4(b). Both figures, whether over VGG or ResNet18, show a roughly linear increase in overhead as $\gamma$ increases. The case where $\gamma=\Gamma_{p}$ represents the scenario where $\gamma_{q},\forall q$ is randomly selected as an integer from the range $[1,3]$, with the average $\gamma_{q}$ being $1.9$ across our experiments. This scenario incurs slightly higher overhead due to the variability in $\gamma_{q}$ compared to a uniform expansion factor. Since CIFAR-100 and CIFAR-10 are similar in size ($\sim 177$ MB), their computational and transmission overheads are nearly identical. However, using CIFAR-10 and MNIST in scenarios with different client tasks significantly increases training and transmission times.
Takeaways. SplitWiper/SplitWiper+ eliminate overhead increases as epochs grow through one-way-one-off propagation, ensuring that unrelated clients remain unaffected during unlearning. Despite label expansion overhead, SplitWiper+ reduces computational and communication costs by as much as SplitWiper, achieving over 97-99% reduction in unlearning compared to existing SL methods.
@@ 5.3 Unlearning Effectiveness and Accuracy
The right sides of Table 5 and Table 3 present the accuracy of the label to be unlearned, $Y_{u}$, and other labels, $Y_{o}$, under both learning and unlearning scenarios across the test set. It is evident that all frameworks, including existing ones and ours with varies $\gamma$ and $\varepsilon$, effectively meet G2 (effective unlearning with retained utility) by reducing the accuracy of $Y_{u}$ to zero through unlearning, while simultaneously ensuring that the accuracy of $Y_{o}$ remains unaffected. Considering the CIFAR-100 dataset features a larger number of labels and fewer training examples per label, it poses a higher learning challenge compared to CIFAR-10, resulting in generally lower accuracy across all frameworks on CIFAR-100. Additionally, accuracy on CIFAR-100 is lower with the ResNet18 architecture compared to VGG, as VGG demonstrates better suitability for this specific task in our experiments.
@@
Compared to Vanilla and U-shaped SL, SplitWiper shows a notable accuracy increase for both $Y_{u}$ and $Y_{o}$ in the learning scenario. It surpasses Vanilla SL by 4.77% in CIFAR-10, 4.37% in CIFAR-100, and 2.98% in scenarios involving CIFAR-10 and MNIST. Against U-shaped SL, SplitWiper achieves accuracy improvements of 4.18% in CIFAR-10, 3.82% in CIFAR-100, and 3.54% in mixed CIFAR-10 and MNIST scenarios. The key feature of SplitWiper is freezing client parameters after pre-training, keeping them unaffected during server-side training. When all clients share a server model $\mathcal{F}_{o}^{s}$, back-propagation from $\mathcal{F}_{o}^{s}$ can cause overfitting or hinder learning in client-$k$’s model $\mathcal{F}_{o}^{k}$. Freezing client parameters in SplitWiper enhances efficiency, reduces overhead, and minimizes cross-client interference, improving the accuracy of $Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s})$.
Comparing SplitWiper and SplitWiper+, Figs. 4(c) and 4(d) show the impact of $\varepsilon$ and $\gamma$ on accuracy. As $\gamma$ increases, accuracy decreases slightly but remains within acceptable limits. In SplitWiper+ with $\gamma>1$, higher $\varepsilon$ reduces noise and improves test accuracy. SplitWiper+ with $\varepsilon=0.2$ can even outperform SplitWiper ($\gamma=1$) due to minor noise and label expansion acting as data augmentation [30], enhancing accuracy and generalization. Specifically, SplitWiper+ with $\varepsilon=0.2$ and $\gamma=\Gamma_{p}$ shows an average accuracy increase of 0.37% in CIFAR-10 and 0.12% in CIFAR-100, particularly for $\mathbb{Y}_{o}$, while maintaining comparable performance on $\mathbb{Y}_{u}$. This demonstrates SplitWiper+’s ability to enhance privacy protection without compromising utility.
Takeaways. SplitWiper and SplitWiper+ effectively unlearn upon request while preserving the accuracy of retained data. The unlearning process in SplitWiper is over 108% more effective than existing SLs, with SplitWiper+ further enhancing accuracy by an additional 0.37%.

### v4

@@ 5.2 Masked Label Sharing in SplitWiper+
Building on the proposed label expansion strategy, we examine how SplitWiper+ implements masked label sharing to enhance privacy, distinguishing it from the SplitWiper framework.
Masked label sharing. In SplitWiper+, after freezing the pre-trained model, clients employ an expansion-and-masking scheme before sharing their data with the server. In contrast, SplitWiper involves clients directly transmitting the obtained data $\{Y_{q},V_{q}^{k}\}$ to the server after freezing the pre-trained model, without performing any additional masking step. The transformation of $\{Y_{q},V_{q}^{k}\}$ into $\{Y_{e}^{*},{V_{e}^{*}}^{k}\}$ enables clients to transmit masked labels and their corresponding DP-protected intermediate values to the server, ensuring that the original information remains concealed.
@@
Consider that $\mathbb{V}_{exp}=\{{V_{1}^{*}},\cdots,{V_{E}^{*}}\}$ is expanded by $\mathbb{V}=\{{V}_{1},\cdots,{V}_{Q}\}$ using a given method $\mathcal{G}_{V}$, where $\forall{V}_{q}\in\mathbb{V}$, $\gamma_{q}$ copies with $\varepsilon$-DP-based noise are generated as $\{{V_{a_{1}}^{*}},\cdots,{V_{a_{\gamma_{s}}}^{*}}\}\in\mathbb{V}_{exp}$. The probability that any two elements in $\mathbb{V}_{exp}$ correspond to the same element in $\mathbb{V}$ satisfies $2\varepsilon$-DP.
@@
We introduce the label demasking on the client side when a client needs to utilize the complete model, i.e., $Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s})$. This step converts the predicted masked labels received from the server back into their original.
Given that $\mathcal{G}_{Y}$ is a one-to-many label expansion method, its inverse, $\mathcal{G}_{Y}^{-1}$, operates as a many-to-one function, defined as:
@@
Each masked label corresponds to a real label, and since the mapping $\mathcal{G}_{Y}$ is shared among all clients, they can easily demask the server’s predicted results using the inverse mapping $\mathcal{G}_{Y}^{-1}$. This demasking process ensures the integrity and applicability of the model’s predictions. The demasked labels are then used to compute the performance $\mathcal{T}(Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s}),D_{o}^{k})$, ensuring evaluation is based on accurate and relevant data, reflective of true model performance in real-world settings.
Label protection in unlearning process. When the initiating client $k$ requests unlearning, only client $k$ participates in the subsequent update process. Other clients do not receive any additional updated information. The server, upon receiving the updated masked labels and corresponding DP-protected intermediate values shared by client $k$, gains insight into the aspects of the unlearning process. However, since the labels are masked, the server cannot accurately infer the quantity or category of the real labels involved in the unlearning from the changes in the number of masked labels. Thus, we deem that the privacy associated with the unlearning labels remains protected, effectively safeguarding sensitive information from inadvertent disclosure.
@@
    Training on private datasets. When training on private datasets, an external attacker with partial access to a victim’s data cannot construct a complete shadow dataset. While [34] assumes the attacker knows all labels in the victim’s dataset, this is unrealistic unless using a well-known public dataset. Balancing positive and negative cases and diversifying shadow models is crucial for generalization in the attack model. Without full label knowledge, inferring the complete input-output mapping is challenging, restricting the attacker’s ability to replicate the training environment and conduct effective inference.
@@
Compared to SplitWiper, label expansion in SplitWiper+ introduces additional computational overhead and increases the size of the data transmitted from clients to the server, thereby increasing communication costs. Thanks to the one-way-one-off propagation scheme, these additional overheads are incurred only once, preserving the significant advantages over existing SL methods, particularly in client-side efficiency. A detailed analysis of client-side and server-side overhead complexities are shown in Appendix B, where Table 5 and Table 6 respectively present a comparison of the client-side and server-side overhead among SplitWiper, SplitWiper+, and existing Vanilla and U-shaped SL frameworks.
@@ 6.1 Experimental Settings
Our experiments run on NVIDIA PCIe A100, 2 $\times$ 40GB. We use Python 3.7.16 and PyTorch 2.3.1 to build and train our framework.
We evaluate the effectiveness, efficiency, and privacy level of SplitWiper and SplitWiper+ with $K=5$ clients and a single server, where client data is distributed in a non-IID manner.
@@
SplitWiper and SplitWiper+ are versatile, accommodating a variety of unlearning scenarios including class-, client-, and sample-level tasks. The experiments specifically focus on class-level unlearning tasks for clarity of presentation [40]. Further experiments on sample-level unlearning tasks are presented in Appendix D. In our settings, $\mathbb{C}_{u}$ selects a subset of label, denoted as $\mathbb{Y}_{u}$ for unlearning, aiming to effectively erase all related information without affecting other clients $\mathbb{C}_{o}$, or other labels $\mathbb{Y}_{o}$. We set $\lvert\mathbb{Y}_{u}\rvert=1$, i.e., only a specific label is required for unlearning. The unlearned label set, $\mathbb{Y}_{u}$, is exclusively possessed by $\mathbb{C}_{u}$, aligned with the non-IID data distribution setting and enabling a clearer demonstration of unlearning effectiveness.
@@
    Clients with different tasks. Clients utilize different datasets for separate tasks, with distinct labels among the clients.
@@
For experiments where clients perform the same task, we use the CIFAR-10 [41, 42] and CIFAR-100 [43, 44] datasets, both widely utilized in the fields of computer vision and neural networks. CIFAR-10 consists of $60,000$ $32\times 32$ color images distributed across $10$ classes, while CIFAR-100 includes $100$ classes with $600$ images per class. In the CIFAR-10 experiments, the designated label $Y_{u}$ is set to “bird,” and in CIFAR-100, it is set to “aquarium-fish.” For experiments where clients perform different tasks, $3$ clients utilize the CIFAR-10 dataset, while $2$ clients use the MNIST dataset [45], which comprises $60,000$ $28\times 28$ grayscale images of handwritten digits. In this setup, the label “bird” from CIFAR-10 remains designated as $Y_{u}$.
@@
In scenarios where clients train on the same task, each client’s model, $\mathcal{F}_{o}^{k},\forall k$, adheres to a uniform structure with the same input and output sizes across all clients. Conversely, in scenarios with clients engaged in different tasks, the input size of each $\mathcal{F}_{o}^{k}$ is tailored based on the specific training dataset $D_{o}^{k}$ used locally. However, the output size remains consistent to meet the input requirements of the server model $\mathcal{F}_{o}^{s}$. This design ensures that despite the diverse data types and tasks, all client models can interface correctly with the centralized server model within SL.
@@
Benchmark and experimental groups. We use non-SISA unlearning on Vanilla and U-shaped SL [48] as benchmarks to evaluate SplitWiper. To ensure fairness, both benchmarks are initialized with pre-trained results from SplitWiper. The client and server components in these frameworks are trained for the same $M$ epochs as SplitWiper during learning and retraining for unlearning requests. By standardizing initial conditions and training durations, we systematically compare the efficiency and effectiveness of SplitWiper against the benchmarks.
@@
To assess the impact of various label expansion strategies in SplitWiper and SplitWiper+, we categorize experiments based on different expansion factors. Specifically, we define four experimental groups: three groups with uniform label expansions of $\gamma=1$, $2$, or $3$, and one group with a variable expansion factor, $\gamma=\Gamma_{p}$. For $\gamma=1$ (SplitWiper), labels are shuffled and re-indexed without altering their actual count, similar to the Vanilla SL approach where real labels are shared. For $\gamma=2$ or $\gamma=3$ (SplitWiper+), clients expand labels to $\gamma$ instances and apply differential privacy ($\varepsilon$-DP) with $\varepsilon=0.1$ or $0.2$ to protect intermediate values.
In the variable scenario ($\gamma=\Gamma_{p}$), the expansion factor $\gamma_{q}$ for each label $Y_{q}$ is randomly selected from a predefined range, $\Gamma_{p}=[1,3]$. This configuration enables us to examine the effects of variability in the expansion factor on both label privacy and system performance.
Performance metrics. G1 represents a significant advancement by enabling absent clients and involving only those requesting unlearning, laying the groundwork for SISA in SL. While not easily assessed quantitatively, G1 serves as the foundational experiment setting and premise for achieving G2 and G3. We analyze the unlearning effectiveness and efficiency of existing SL frameworks, SplitWiper, and SplitWiper+. We measure label accuracy during learning and unlearning phases for effectiveness, and track training and transmission times for efficiency. We focus on client-side overhead and use a 100 Mbps bandwidth setting to simulate realistic conditions. This approach demonstrates the benefits of the one-way-one-off propagation.
To further evaluate G3, we assess the robustness of our privacy-preserving method. We examine the server’s ability to infer real labels from masked labels and intermediate outputs. In SplitWiper+, the server uses clustering algorithms on received data pairs to group similar labels. The metric is the probability of correctly matching masked labels to real ones.
@@ 6.2 Unlearning Efficiency and Overhead
The results highlight the efficacy of SplitWiper in achieving G1 (independent unlearning). Training and transmission times in SplitWiper are significantly lower than those in Vanilla and U-shaped SL across all data distributions and scenarios. Crucially, during unlearning, only the requesting clients, $\mathbb{C}_{u}$, incur additional costs, while other clients, $\mathbb{C}_{o}$, remain unaffected, as the server uses stored data from the learning phase to handle unlearning, minimizing overall client-side expenses and disruptions. The time overhead for the label expansion strategy discussion between clients (cf. Algorithm 3) is negligible, measured in milliseconds, and does not significantly impact our analysis of training and transmission times.
In SplitWiper and SplitWiper+, clients only need to train the pre-trained model over local data once. They then mask labels according to the label expansion strategy and apply DP protection to the corresponding intermediate values before transmitting them to the server in a one-way propagation manner. Subsequent training phases focus exclusively on the server, disconnecting from client-side computations. Therefore, the training and transmission times for clients are fixed values dependent on $\gamma$ and remain unaffected by the number of training epochs $M$. In contrast, in Vanilla and U-shaped SL, where each epoch involves client participation, the computational and transmission times increase linearly with $M$, as listed in Table 5 (cf. Appendix B).
@@
To illustrate the impact of the label expansion factor $\gamma$ on client-side overhead in SplitWiper+, we plot the average training and transmission times for all clients across various $\gamma$ values in Fig.5(a) and Fig.5(b). Both figures, for VGG and ResNet18, show a roughly linear increase in overhead as $\gamma$ grows. The case $\gamma=\Gamma_{p}$ represents scenarios where $\gamma_{q}$ is randomly selected as an integer from $[1,3]$, with an average $\gamma_{q}$ of $1.9$. This introduces slightly higher overhead due to variability in $\gamma_{q}$ compared to a uniform expansion factor. Since CIFAR-100 and CIFAR-10 are similar in size ($\sim 177$ MB), their computational and transmission overheads are nearly identical. However, using CIFAR-10 and MNIST in scenarios with differing client tasks significantly increases training and transmission times.
Takeaway (faster). SplitWiper/SplitWiper+ mitigate overhead growth with one-way-one-off propagation, ensuring unrelated clients remain unaffected during unlearning. While SplitWiper+ introduces label expansion overhead, it matches SplitWiper in reducing computational and communication costs, achieving a 97-99% reduction in unlearning overhead compared to existing SLs.
@@ 6.3 Unlearning Effectiveness and Accuracy
The right sides of Tables 123 show the accuracy of the labels to be unlearned ($Y_{u}$) and other labels ($Y_{o}$) under both learning and unlearning scenarios across the test set. All frameworks, including existing ones and ours with varying $\gamma$ and $\varepsilon$, effectively satisfy G2 (effective unlearning with retained utility). They achieve zero accuracy for $Y_{u}$ during unlearning (consistent with retraining results [40, 49, 50]) while maintaining the accuracy of $Y_{o}$. CIFAR-100, with more labels and fewer training examples per label, presents a greater learning challenge than CIFAR-10, resulting in lower accuracy across all frameworks. Additionally, ResNet18 performs worse than VGG on CIFAR-100, as VGG proves better suited for this task in our experiments.
@@
Compared to Vanilla and U-shaped SL, SplitWiper achieves notable accuracy improvements for both $Y_{u}$ and $Y_{o}$ in the learning scenario. It surpasses Vanilla SL by 4.77% on CIFAR-10, 4.37% on CIFAR-100, and 2.98% in CIFAR-10 and MNIST scenarios. Against U-shaped SL, it shows gains of 4.18%, 3.82%, and 3.54% in the same respective scenarios. SplitWiper’s key innovation is freezing client parameters after pre-training, preventing them from being influenced during server-side training. Shared server model back-propagation ($\mathcal{F}_{o}^{s}$) can cause overfitting or hinder learning in a client’s model ($\mathcal{F}_{o}^{k}$). Freezing client parameters enhances efficiency, reduces overhead, and minimizes cross-client interference, improving the accuracy of $Concat(\mathcal{F}_{o}^{k},\mathcal{F}_{o}^{s})$.
Figs.5(c) and 5(d) compare SplitWiper and SplitWiper+, showing the effects of $\varepsilon$ and $\gamma$ on accuracy. While increasing $\gamma$ slightly reduces accuracy, it remains acceptable. In SplitWiper+ with $\gamma>1$, higher $\varepsilon$ reduces noise and improves test accuracy. With $\varepsilon=0.2$, SplitWiper+ can even outperform SplitWiper ($\gamma=1$) due to minimal noise and label expansion functioning as data augmentation[51, 52], enhancing accuracy and generalization. Specifically, SplitWiper+ with $\varepsilon=0.2$ and $\gamma=\Gamma_{p}$ achieves an average accuracy gain of 0.37% on CIFAR-10 and 0.12% on CIFAR-100, particularly improving $\mathbb{Y}_{o}$ while maintaining comparable performance on $\mathbb{Y}_{u}$. This highlights SplitWiper+’s ability to strengthen privacy protection without compromising utility.
Takeaway (more effective). SplitWiper/SplitWiper+ effectively unlearn upon request while preserving the accuracy of retained data. The unlearning process in SplitWiper is over 8% more effective than existing SL, with SplitWiper+ further enhancing accuracy by an additional 0.37%.